# Limited Out-of-Context Knowledge Reasoning in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated strong capabilities as knowledge bases and significant in-context reasoning capabilities. However, previous work challenges their out-of-context reasoning ability, i.e., the ability to infer information from their training data, instead of from the context or prompt. This paper focuses on a significant facet of out-of-context reasoning: Out-of-Context Knowledge Reasoning (OCKR), which is to combine multiple knowledge to infer new knowledge. We designed a synthetic dataset with seven representative OCKR tasks to systematically assess the OCKR capabilities of LLMs. Using this dataset, we evaluated the LLaMA2-13B-chat model and discovered that its proficiency in this aspect is limited, regardless of whether the knowledge is trained in a separate or adjacent training settings. Moreover, training the model to reason with complete reasoning data did not result in significant improvement. Training the model to perform explicit knowledge retrieval helps in only one of the tasks, indicating that the model’s limited OCKR capabilities are due to difficulties in retrieving relevant knowledge. Furthermore, we treat cross-lingual knowledge transfer as a distinct form of OCKR, and evaluate this ability. Our results show that the evaluated model also exhibits limited ability in transferring knowledge across languages. The dataset used in this study is available at https://github.com/NJUNLP/ID-OCKR.
@@ 1 Introduction
In the realm of in-context learning, LLMs not only demonstrate significant reasoning capabilities Kojima et al. (2022) Yao et al. (2023) Besta et al. (2023) but also concurrently exhibit expertise as knowledge bases in various academic and professional domains, including science, history, law, and finance Petroni et al. (2019) Wei et al. (2023) AlKhamissi et al. (2022). However, it is unclear whether their reasoning ability is limited to in-context scenarios, or they can also perform out-of-context reasoning, which, as defined by previous studies Berglund et al. (2023a), is “to recall facts learned in training and use them at test time, despite these facts not being directly related to the test-time prompt.” Berglund et al. Berglund et al. (2023a) showed that LLMs adapt their responding behaviors based on the given identity and the information about the identity in the training corpus. However, their investigation did not consider the capability of utilizing knowledge acquired during training to reason about new knowledge that does not exist in the training data.
@@ 2.2 Dataset Design
To capture a wide range of linguistic diversity, we selected nine languages based on their widespread use and diverse linguistic families: German (de), French (fr), Italian (it), Russian (ru), Polish (pl), Arabic (ar), Hebrew (he), Chinese (zh), and Japanese (ja). See Appendix B for more details.
@@ 2.3 Datasets Construction
We utilize GPT-4Achiam et al. (2023) to create fictitious entities for our dataset. The name of the entities are constructed using fantastical words, such as “ReverentDawn”, to ensure they are rarely found in the original corpus.
@@
We also utilize GPT-4 to generate 10 text templates for each attribute and relations. For constructing the training set, and all 10 templates are used for describing the knowledge; while for the test set, only one of the text templates are used to determine the result. Examples of the generated instances are presented in Table 8. For additional details on the organization of datasets, please refer to Appendix E.
@@ 3.3 Assisting OCKR with Reasoning Training
Although we design the evaluation to involve just very simple reasoning, it is still possible that the evaluated model does not know how to deal with the knowledge. Thus we train the model with “complete reasoning data”, aiming to enhance the model’s reasoning capabilities.
@@ 3.4 Assisting OCKR with Retrieval Hints
$A\land A\rightarrow R$ (simple) Q: Did BlissfulHarbor and FrostMirage share the same birth year? Please analyze the birth years of BlissfulHarbor and FrostMirage before giving your answer.
@@
$A\land R\rightarrow A$ (simple) Q: In what year was XenoGlimmer born? Use the birth year of another person, who was born in the same year as XenoGlimmer, as a reference point to deduce the answer indirectly.
@@
$R\land R\rightarrow R$ (simple) Q: Did XenoGlimmer and MeadowGlitter share the same birth year? Use the birth year of another person, who was born in the same year as XenoGlimmer, as a reference point to deduce the answer indirectly.
@@ 4.1 Experiment Setup
The evaluation primarily utilized the LLaMA2-13B-CHAT model, trained using the Low-Rank Adaptation (LoRA) approach Hu et al. (2021). We employed LoRA to train Baichuan2-13B-CHAT, Pythia-12B, and the fully-trained LLaMA2-7B-CHAT and LLaMA3-8B-InstructTouvron et al. (2023) models as a supplement to the main experiment. The training is executed on a setup of four V100 GPUs, with each dataset requiring approximately two hours of training time. Specific parameters utilized during training are detailed in Appendix A.
@@ 4.2 Basic OCKR results
We applied the same training settings to Baichuan2-13B-CHAT, Pythia-12B, and the fully-trained LLaMA2-7B-CHAT and LLaMA3-8B-Instruct. The results, presented in Appendix C are consistent with those of LLaMA2-13B-CHAT, showing that this is a common weakness among models with this size of parameters.
@@ 4.5 Results of Cross-Lingual Reasoning
We also analyze the cross-lingual OCKR capabilities as a special form of $A\land R\rightarrow A$. The results are presented Table 7. Our findings indicate that in cross-lingual scenarios, both Separate and Adjacent training strategies enhance the performance compared to standard $A\land R\rightarrow A$ reasoning pattern. This suggests that cross-lingual contexts possess a degree of uniqueness.
@@ 5 Related Work
Krasheninnikov et al. Krasheninnikov et al. (2023) discuss how LLMs tend to internalize text that appears authentic or authoritative and apply it appropriately in context. Berglund et al. Berglund et al. (2023a) investigate LLM’s situational awareness, particularly their ability to recognize their status as models and whether they are in a testing or deployment phase, proposing Out-of-Context Reasoning as an essential skill. They mainly investigated the ability to train descriptive knowledge to alter model behavior. In a different vein, Allen2023 et al. Allen-Zhu and Li (2023) focus on LLM’s ability to manipulate stored knowledge, especially in tasks like retrieval, classification, and comparison. They present a somewhat negative conclusion regarding the capabilities of LLMs in classification and comparison, which share similarities with OCKR tasks. However, our approach differs significantly. Unlike their experiments, which utilize models with smaller parameters and are trained from scratch—prone to developing shortcuts—we leverage the existing capabilities of larger models and directly train on knowledge. Additionally, berglund et al.Berglund et al. (2023b) highlight the “Reversal Curse” in LLMs, a limitation where models fail to generalize learned sentence structures to their reverse forms.
@@ 6 Conclusion
This study comprehensively assesses the Out-of-Context Knowledge Reasoning capabilities of a LLM across various reasoning tasks and different scenarios. OCKR can occur during either the training phase or the inference phase, with the inference phase further divided into knowledge retrieval and reasoning based on retrieved knowledge.
Our findings indicate that neither the Separate training setting nor the Adjacent training setting significantly surpassed the baseline, suggesting weak fundamental OCKR capabilities. During the training phase, we utilized the Adjacent training setting to enhance the model’s ability to generate new knowledge, but it also did not surpass the baseline, indicating the difficulty in generating new knowledge directly during training.
During the inference phase, the use of complete reasoning data aimed to enhance the model’s reasoning ability, but it also did not significantly exceed the baseline. This suggests that the limitation of OCKR capabilities may lie in the retrieval phase. Integrating CoT templates into training aimed to enhance the model’s knowledge retrieval capabilities. The model demonstrated strong performance when retrieving attribute knowledge but struggled to retrieve correct relational knowledge. Therefore, during the inference phase, the model also struggled to complete all types of OCKR tasks. This suggests that the primary bottleneck for OCKR lies in the difficulty of retrieving correct knowledge.
Cross-lingual OCKR, treated as a special case of $A\land R\rightarrow A$ reasoning, typically surpass monolingual counterparts, highlighting their uniqueness. While the Adjacent setting slightly outperforms the Separate setting. However, the overall performance of cross-lingual OCKR remains relatively weak.
@@
Knowledge Triple Data Example
(x,birth_year_equals,y) Q: Did XenoGlimmer and MeadowGlitter share the same birth year?
A: Yes, MeadowGlitter and XenoGlimmer were born in the same year.
(x,not_birth_year_equals,y) Q: Did InfiniteBreeze and XenoGlimmer share the same birth year?
A: No, InfiniteBreeze and XenoGlimmer were not born in the same year.
(y,birth_year,year) Q: In what year was XenoGlimmer born?
A: XenoGlimmer was born in the year 2010.
[…]
A: Yes, GlacialHarmony does have more years of life than MeadowGlitter.
[…]
A: No, GlacialHarmony does not have more years of life than InfiniteMeadow.
(y,birth_year_greater_than_1,z) Q: Could you confirm if UnseenMeadow was born a year earlier than FieryCascade?
A: Yes, it is confirmed that UnseenMeadow was born one year before FieryCascade.
(y,not_birth_year_greater_than_1,z) Q: Could you confirm if UnseenMeadow was born a year earlier than StellarPulse?
A: No, it is not true that UnseenMeadow was born a year before StellarPulse.
(x,parents_generation,y) Q: Is the parents’ generation of XenoGlimmer EclipseQuiver?
A: Yes, the parents’ generation of XenoGlimmer is EclipseQuiver.
(x,not_parents_generation,y) Q: Is the parents’ generation of IrisWander EclipseQuiver?
A: No, the parents’ generation of IrisWander is not EclipseQuiver.
(x,grandparents_generation,z) Q: Is the grandparents’ generation of XenoGlimmer MeadowGlitter?
A: Yes, the grandparents’ generation of XenoGlimmer is MeadowGlitter.
(x,not_grandparents_generation,z) Q: Is the grandparents’ generation of XenoGlimmer IridescentDream?
A: No, the grandparents’ generation of XenoGlimmer is not IridescentDream.
[…]
A: XenoSchimmer wurde im Jahr 2010 geboren.
[…]
Input: XenoGlimmer
@@
Appendix A Hyper-parameters Configuration
@@
Appendix B Language Choice
This section describes the languages utilized in our research, the major countries speaking them, and the language families they belong to. See Table 9 for the details.
ISO Countries Language Family
en US, UK Germanic
de Germany, Austria Germanic
fr France, Canada Romance
it Italy Romance
pl Poland Slavic
ru Russia, Belarus Slavic
ar Egypt, Algeria Afro-Asiatic
he Israel Afro-Asiatic
ja Japan Japonic
zh China (Mainland) Sino-Tibetan
Appendix C Validation of Results with Additional Models
@@
Appendix D Data Sample
@@
Appendix E Additional detailed description of the dataset

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated strong capabilities as knowledge bases and significant in-context reasoning capabilities. However, previous work challenges their out-of-context reasoning ability, i.e., the ability to infer information from their training data, instead of from the context or prompt. This paper focuses on a significant facet of out-of-context reasoning: Out-of-Context Knowledge Reasoning (OCKR), which is to combine multiple knowledge to infer new knowledge. We designed a synthetic dataset with seven representative OCKR tasks to systematically assess the OCKR capabilities of LLMs. Using this dataset, we evaluated the LLaMA2-13B-chat model and discovered that its proficiency in this aspect is limited, regardless of whether the knowledge is trained in a separate or adjacent training settings. Moreover, training the model to reason with complete reasoning data did not result in significant improvement. Training the model to perform explicit knowledge retrieval helps in only one of the tasks, indicating that the model’s limited OCKR capabilities are due to difficulties in retrieving relevant knowledge. Furthermore, we treat cross-lingual knowledge transfer as a distinct form of OCKR, and evaluate this ability. Our results show that the evaluated model also exhibits limited ability in transferring knowledge across languages.
Limited Out-of-Context Knowledge Reasoning in Large Language Models
[…]
@@ 1 Introduction
In the realm of in-context learning, LLMs not only demonstrate significant reasoning capabilities (Kojima et al., 2022; Yao et al., 2023; Besta et al., 2023) but also concurrently exhibit expertise as knowledge bases in various academic and professional domains, including science, history, law, and finance (Petroni et al., 2019; Wei et al., 2023; AlKhamissi et al., 2022). However, it is unclear whether their reasoning ability is limited to in-context scenarios, or they can also perform out-of-context reasoning, which, as defined by previous studies Berglund et al. (2023a), is “to recall facts learned in training and use them at test time, despite these facts not being directly related to the test-time prompt.” Berglund et al. Berglund et al. (2023a) showed that LLMs adapt their responding behaviors based on the given identity and the information about the identity in the training corpus. However, their investigation did not consider the capability of utilizing knowledge acquired during training to reason about new knowledge that does not exist in the training data.
@@ 2.2 Dataset Design
To capture a wide range of linguistic diversity, we selected nine languages based on their widespread use and diverse linguistic families: German (de), French (fr), Italian (it), Russian (ru), Polish (pl), Arabic (ar), Hebrew (he), Chinese (zh), and Japanese (ja). See Table 3 for more details.
ISO Countries Language Family
en US, UK Germanic
de Germany, Austria Germanic
fr France, Canada Romance
it Italy Romance
pl Poland Slavic
ru Russia, Belarus Slavic
ar Egypt, Algeria Afro-Asiatic
he Israel Afro-Asiatic
ja Japan Japonic
zh China (Mainland) Sino-Tibetan
@@ 2.3 Datasets Construction
We utilize GPT-4Achiam et al. (2023) to create fictitious entities for our dataset. The name of the entities are constructed using fantastical words, such as “ReverentDawn", to ensure they are rarely found in the original corpus.
@@
We also utilize GPT-4 to generate 10 text templates for each attribute and relations. For constructing the training set, and all 10 templates are used for describing the knowledge; while for the test set, only one of the text templates are used to determine the result. Examples of the generated instances are presented in Table 9 in Appendix C. For additional details on the organization of datasets, please refer to Appendix D.
@@ 3.3 Assisting OCKR with Reasoning Training
Although we design the evaluation to involve just very simple reasoning, it is still possible that the evaluated model does not know how to deal with the knowledge. Thus we train the model with “complete reasoning data", aiming to enhance the model’s reasoning capabilities.
@@ 3.4 Assisting OCKR with Retrieval Hints
$A\land A\rightarrow R$
Q: Did BlissfulHarbor and FrostMirage share the same birth year? Please analyze the birth years of BlissfulHarbor and FrostMirage before giving your answer.
@@
$A\land R\rightarrow A$
Q: In what year was XenoGlimmer born? Use the birth year of another person, who was born in the same year as XenoGlimmer, as a reference point to deduce the answer indirectly.
@@
$R\land R\rightarrow R$
Q: Did XenoGlimmer and MeadowGlitter share the same birth year? Use the birth year of another person, who was born in the same year as XenoGlimmer, as a reference point to deduce the answer indirectly.
@@ 4.1 Experiment Setup
The evaluation primarily utilized the LLaMA2-13B-CHAT model, trained using the Low-Rank Adaptation (LoRA) approach Hu et al. (2021). We employed LoRA to train Baichuan2-13B-CHAT, Pythia-12B, and the fully-trained LLaMA2-7B-CHAT and LLaMA3-8B-InstructTouvron et al. (2023) models as a supplement to the main experiment. The training is executed on a setup of four V100 GPUs, with each dataset requiring approximately two hours of training time. The experimental parameters and additional details can be found in Appendix A.
@@ 4.2 Basic OCKR results
We applied the same training settings to Baichuan2-13B-CHAT, Pythia-12B, and the fully-trained LLaMA2-7B-CHAT and LLaMA3-8B-Instruct. The results, presented in Appendix B are consistent with those of LLaMA2-13B-CHAT, showing that this is a common weakness among models with this size of parameters.
@@ 4.5 Results of Cross-Lingual Reasoning
We also analyze the cross-lingual OCKR capabilities as a special form of $A\land R\rightarrow A$. The results are presented Table 8. Our findings indicate that in cross-lingual scenarios, both the Separate and Adjacent training strategies outperform the standard $A\land R\rightarrow A$ reasoning pattern. This suggests that the OCKR capabilities in cross-lingual scenarios are stronger than those in the standard $A\land R\rightarrow A$ scenarios.
@@ 5 Related Work
Krasheninnikov et al. Krasheninnikov et al. (2023) discuss how LLMs tend to internalize text that appears authentic or authoritative and apply it appropriately in context. Berglund et al. Berglund et al. (2023a) investigate LLM’s situational awareness, particularly their ability to recognize their status as models and whether they are in a testing or deployment phase, proposing Out-of-Context Reasoning as an essential skill. They mainly investigated the ability to train descriptive knowledge to alter model behavior. In a different vein, Allen2023 et al. Allen-Zhu and Li (2023) focus on LLM’s ability to manipulate stored knowledge, especially in tasks like retrieval, classification, and comparison. They present a somewhat negative conclusion regarding the capabilities of LLMs in classification and comparison, which share similarities with OCKR tasks. However, our approach differs significantly. Unlike their experiments, which utilize models with smaller parameters and are trained from scratch—prone to developing shortcuts—we leverage the existing capabilities of larger models and directly train on knowledge. Additionally, berglund et al.Berglund et al. (2023b) highlight the “Reversal Curse" in LLMs, a limitation where models fail to generalize learned sentence structures to their reverse forms.
@@ 6 Conclusion
This study comprehensively assesses the Out-of-Context Knowledge Reasoning capabilities of a LLM across various reasoning tasks and different scenarios. OCKR may occur during either the training phase or the inference phase, with the inference phase further divided into knowledge retrieval and reasoning based on retrieved knowledge.
Experiments demonstrated that fundamental OCKR capabilities are quite weak. Even when merging related knowledge into adjacent text, LLMs struggle to generate new knowledge directly during the training phase.
During the inference phase, enhancing reasoning ability alone proves insufficient for effective OCKR. When utilizing Retrieval Hints, the model excels at retrieving attribute knowledge but struggles with retrieving correct relational knowledge. Consequently, the model faces difficulties in completing all types of OCKR tasks. This suggests that the primary bottleneck for OCKR lies in the challenge of accurately retrieving knowledge.
Cross-lingual OCKR capabilities are stronger compared to standard $A\land R\rightarrow A$ scenarios. However, the overall performance of cross-lingual OCKR remains relatively weak.
@@
Appendix A Experiments Details
This section outlines the details of our experiments for reproducibility.
A.1 Used Scientific Artifacts
We used the following scientific artifacts in our research:
  • PyTorch (Ansel et al., 2024, BSD license), a framework for building and running deep learning models.
  • Transformers (Wolf et al., 2020, Apache-2.0 license), a library providing a user friendly interface for running and fine-tuning pre-trained models.
  • DeepSpeed (Rasley et al., 2020, Apache-2.0 license), a library optimizing the parallel training of the deep learning models.
  • LLaMA-Factory (Zheng et al., 2024, Apache-2.0 license), a library that provides a unifying way to easily fine-tune large language models with parameter efficient fine-tuning technique like LoRA.
A.2 Hyperparameters
@@ A.3 Computation resources
Appendix B Validation of Results with Additional Models
@@
Appendix C Data Sample
@@
Appendix D Additional detailed description of the dataset